A solution is made by dissolving 2.5 grams ofHCN in 558 ml of water. Calculate the ph solution.
1
Expert's answer
2020-12-03T13:52:48-0500

pKa(HCN) = 9.2, which means Ka(HCN) = 10^(-9.2) = 6.31*10^(-10).

Ka = [H+]*[CN-]/[HCN] ≈ [H+]^2/c(HCN)

c(HCN) = n(HCN)/V(solution) = m(HCN)/Mr(HCN)/V(solution) = (2.5 g)/(27.0253 g/mol)/(558 ml) = 0.16578 mol/L

[H+] = (Ka*c(HCN))^1/2 = 1.022*10^(-5)

pH = - log([H+]) = 4.99 ≈ 5